# LPENSL-TH-01/24

### v2

@@ 2 The reflection algebra for the XYZ open spin chain
In these definitions, $\xi_{1},\ldots,\xi_{N}$ are arbitrary inhomogeneity parameters. The boundary monodromy matrix $\mathcal{U}_{-}(\lambda)$, as well as the bulk monodromy matrix $T(\lambda)$, act on the tensor product $\mathcal{H}_{0}\otimes\mathcal{H}$, where $\mathcal{H}_{0}\simeq\mathbb{C}^{2}$ is the so-called auxiliary space, whereas $\mathcal{H}=\otimes_{n=1}^{N}\mathcal{H}_{n}$, with $\mathcal{H}_{n}\simeq\mathbb{C}^{2}$, is the quantum space of the model. In other words, $\mathcal{U}_{-}(\lambda)$ and $T(\lambda)$ are to be understood as $2\times 2$ matrices (as written in the auxiliary space) with operators entries in $\mathcal{H}$, and the products in (2.14), (2.15) and (2.16) are to be understood as matrix multiplication in $\mathcal{H}_{0}$. In particular, we shall denote
@@ 3.1 Transfer matrix properties
The transfer matrix $\mathcal{T}(\lambda)$ is an even elliptic polynomial [^4] We say that a function $f$ is an even elliptic polynomial of order (or degree) $2n$ if it is an even function which in addition satisfies the quasi-periodicity properties: $f(\lambda+\pi)=f(\lambda),\qquad f(\lambda+\pi\omega)=\left(-e^{-2i\lambda-i\pi\omega}\right)^{2n}f(\lambda).$ (3.3) It means that there exist constants $C,\lambda_{1},\ldots,\lambda_{n}$ such that $f(\lambda)=C\prod_{j=1}^{n}\operatorname{\theta_{\mathrm{s}}}(\lambda,\lambda_{j}).$ (3.4) in $\lambda$ of order $62N+$ 6, i.e. it satisfies the following parity and quasi-periodicity properties in $\lambda$ w.r.t. the periods $\pi$ and $\pi\omega$:
@@
$\mathcal{T}(\xi_{0}^{(0)})=\tau_{0}\,\mathbb{I}$ with $\text{with }\,\tau_{0}=(-1)^{N}\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,\det_{q}T(0),$ (3.7)
$\mathcal{T}(\xi_{-1}^{(0)})=\tau_{-1}\,\mathbb{I}$ with $\text{with }\,\tau_{-1}=c^{z}_{-}\,c^{z}_{+}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,\det_{q}T\Big{(}\frac{\pi}{2}\Big{)},$ (3.8)
$\mathcal{T}(\xi_{-2}^{(0)})=\tau_{-2}\,\mathbb{I}$ with $\text{with }\,\tau_{-2}=-e^{i\big{(}2\sum_{n=1}^{N}\xi_{n}-(N+3)\eta-\frac{3\pi\omega}{2}\big{)}}\,c^{x}_{-}\,c_{+}^{x}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,\det_{q}T\Big{(}\!-\!\frac{\pi\omega}{2}\Big{)},$ (3.9)
$\mathcal{T}(\xi_{-3}^{(0)})=\tau_{-3}\,\mathbb{I}$ with $\text{with }\,\tau_{-3}=e^{i\big{(}2\sum_{n=1}^{N}\xi_{n}-(N+3)\eta-\frac{3\pi\omega}{2}\big{)}}\,c^{y}_{-}\,c_{+}^{y}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,\det_{q}T\Big{(}\!-\!\frac{\pi+\pi\omega}{2}\Big{)}.$ (3.10)
@@ 3.2 SoV basis
Then, one can show similarly as in [1] that, for almost any choice of the co-vector $\langle\,S\,|$ and of the inhomogeneity parameters satisfying (3.15), the states
@@ 3.4 Characterisation in terms of a functional $TQ$-equation
The proof of this proposition is standard, we have mainly to observe that, from the $TQ$-equation (3.37), it follows that $\tau(\lambda)$ is an even elliptic polynomial in $\lambda$ of order $62N+$ 6. Moreover, the identities
@@
Such a description of the transfer matrix spectrum and eigenstates in terms of even elliptic polynomial $Q$-functions solution of a homogeneous $TQ$-equation (3.37) in a given sector $M<N$ is of course not complete. A complete description of the general open XYZ transfer matrix spectrum in terms of $Q$-functions solution of a homogeneous [^5] It is also possible to have, in general XYZ open spin chains, a formulation of the spectrum in terms of elliptic polynomial solutions of a $TQ$-equation with an additional inhomogeneous term, as proposed in [103] (see also [102] where another inhomogeneous formulation was derived - and proved to be complete - from SoV in the general open XXZ case). Besides the presence of the inhomogeneous term, the $TQ$-equation of [103] also differs from (3.36)-(3.37) by the fact that it involves a couple of non-even related $Q$-functions, instead of a single even one. We do not consider this kind of formulation in our present paper. $TQ$-equation is still unknown. Note however that, if the constraint (3.35) is satisfied with some choice of $M$ and $\boldsymbol{\varepsilon}$, it is also satisfied, with the same boundary parameters, for $M^{\prime}=N-1-M$ and $\boldsymbol{\varepsilon^{\prime}}=-\boldsymbol{\varepsilon}$. It is therefore quite possible that, as conjectured in [94, 95] in the XXZ case and in [96] in the XYZ case, the solutions $Q(\lambda)$ (3.36) of degree $2M$ of (3.37) with $\boldsymbol{\varepsilon}$, together with the solutions of degree $2M^{\prime}$ with $M^{\prime}=N-1-M$ and $\boldsymbol{\varepsilon^{\prime}}=-\boldsymbol{\varepsilon}$, provide the complete spectrum and eigenstates of the transfer matrix, at least for some range of the boundary parameters.
@@ 4 Scalar products of separate states: main results
for some $\boldsymbol{\varepsilon},\boldsymbol{\varepsilon^{\prime}}\in\{\pm 1\}^{6}$. Here ${\textsc{a}}_{\boldsymbol{\varepsilon}}$ denotes the solution of (3.29) given by (3.32), whereas ${}_{\boldsymbol{\varepsilon}}\langle\,\mathbf{h}\,|$ and $|\,\mathbf{h}\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}$, $\mathbf{h}\equiv(h_{1},\ldots,h_{N})\in\{0,1\}^{N}$, stand for the states (3.17) and (3.20) with function $\mathbf{A}$ (3.16) given by $\mathbf{A}_{\boldsymbol{\varepsilon}}$ and $\mathbf{A}_{\boldsymbol{\varepsilon^{\prime}}}$ respectively666 More generally, the states ${}_{\boldsymbol{\varepsilon}}\langle\,\mathbf{h}\,|$ and $|\,\mathbf{h}\,\rangle_{\boldsymbol{\varepsilon}}$ in (4.1) and (4.4) may stand for any SoV basis elements diagonalising the transfer matrix $\mathcal{T}(\lambda)$ as in Theorem 3.1, provided that it satisfies (3.21) for a given normalisation constant $\mathcal{N}_{0}$, and that it also satisfies (4.3). For instance, such a basis could also be constructed as in [73] by using the Vertex-IRF transformation.. Note that we have
@@ 4.1 The original SoV determinant representation
${}_{\boldsymbol{\varepsilon}}\langle\,P\,|\,Q\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}={}_{\boldsymbol{\varepsilon}}\langle\,Q\,|\,P\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}$
$=P(\xi_{n}^{(0)})\,Q(\xi_{n}^{(0)})\!\sum_{\mathbf{h}\in\{0,1\}^{N}}\prod_{n=1}^{N}\left[-\frac{{\textsc{a}}_{\boldsymbol{\varepsilon}}(\frac{\eta}{2}+\xi_{n})}{{\textsc{a}}_{\boldsymbol{\varepsilon^{\prime}}}(\frac{\eta}{2}-\xi_{n})}\frac{P(\xi_{n}^{(1)})\,Q(\xi_{n}^{(1)})}{P(\xi_{n}^{(0)})\,Q(\xi_{n}^{(0)})}\right]^{h_{n}}\frac{V(\xi_{1}^{(1-h_{1})},\ldots,\xi_{N}^{(1-h_{N})})}{V(\xi_{1},\ldots,\xi_{N})}.$ (4.5)
@@
Let $\Xi_{N}$ be the space of even elliptic polynomials of order $2(N-1)$, i.e. of holomorphic functions $f$ on $\mathbb{C}$ satisfying the properties
@@ 4.2 Slavnov’s type determinant representation
×$\qquad\times\bigg{[}\delta_{j,k}\,\frac{\partial}{\partial\mu}\left(\log\frac{\mathbf{A}_{\boldsymbol{\varepsilon}}(\mu)\,Q(\mu-\eta)}{\mathbf{A}_{\boldsymbol{\varepsilon}}(-\mu)\,Q(\mu+\eta)}\right)_{\mu=q_{i}}+K(q_{i}-q_{j})-K(q_{i}+q_{j})\bigg{]},$ (4.17)
@@ 5 Scalar products of separate states: details of the computations
In this section, as well as in the sections 5.2 and 5.3, we consider the case $\boldsymbol{\varepsilon}=\boldsymbol{\varepsilon^{\prime}}$, $P$ and $Q$ being of the same order $2M$. Since $\boldsymbol{\varepsilon}$ is fixed, we shall slightly simplify the notations and denote $\{a_{i}\}_{i=1,\ldots 6}=\{-\epsilon_{i}^{\sigma}\alpha_{i}^{\sigma}\}_{i=1,2,3;\sigma=\pm}$, so that
@@
in the expression (4.9) of the scalar product of ${}_{\boldsymbol{\varepsilon}}\langle\,P\,|$ and $|\,Q\,\rangle_{\boldsymbol{\varepsilon}}$, we obtain [^8] Here and in the following we are using the compact notation $(PQ)(x)=P(x)Q(x)$.
@@
so that it is an odd elliptic function of $\lambda$. Hence, the sum of the residues at its poles in an elementary cell vanishes. Moreover, if $\mathsf{F}_{i,j}$ has a pole at some point $x$, it has also a pole at $-x$, and
@@
Without any particular assumption on the parameters, i.e. keeping $q_{1},\ldots,q_{M}$ generic, it is possible to further transform the above determinant of size $2M$ by using, as in [3, 5], the usual formula for the determinant of block matrices:
@@
To conclude this section, let us briefly comment about the case in which 2 of the 6 boundary parameters are opposite from each others, say for instance $a_{5}=-a_{6}$. We have already mentioned that, in this special case, it was possible to apply the strategy of [4, 5] to compute the scalar products, see the footnote 9. The choice $a_{5}=-a_{6}$ also induces some simplifications in the new approach presented in this section: under this hypothesis, we can indeed simply choose $\gamma=a_{5}$ in (5.12)-(5.1), so that the functions (5.1) no longer have a pole in $\gamma$. This means that the terms of the form $\mathsf{X}_{\gamma}(u)\,\mathsf{M}_{\{a\}}(\gamma,v)$ in the expressions (5.1)-(5.1) vanish. Therefore, in particular, the matrix $\mathcal{G}^{(2,2)}$ (5.1) is simply a diagonal matrix, instead of being the sum of a diagonal matrix and a rank one matrix, so that the explicit expression of its determinant – and hence of the constant $\mathsf{c}_{Q}$ (5.34) – is much simpler. The general formula issued from (5.26), in which we do not suppose $q_{1},\ldots,q_{M}$ to be a solution of the Bethe equations, would also be much less cumbersome in that case. However, this does not induce any particular simplification in our main result (4.14), once we impose the Bethe equations for $q_{1},\ldots,q_{M}$ and normalise adequately the scalar product.
@@
We can now transform the determinant of the matrix $\mathcal{M}\equiv(\mathsf{M}_{\{a^{\prime}\}}(\xi_{i},w_{j}))_{1\leq i,j\leq N}$ according to the previous approach, but we have to adapt slightly this approach since now $2M<N$. We define the following generalisations of the functions (5.1):
@@
The last column in the determinant of the numerator of (5.92) being identically 00, we obtain the result.
@@ 6 Conclusion
The new scalar product formulas that we have obtained here constitute the first step towards the computation of the correlation functions. Building on this result, we intend to address the problem of the generalisation of the results of [90, 91, 92] to the XYZ open spin chain in a forthcoming publication.

### v3

@@ 2 The reflection algebra for the XYZ open spin chain
In these definitions, $\xi_{1},\ldots,\xi_{N}$ are arbitrary inhomogeneity parameters. The boundary monodromy matrix $\mathcal{U}_{-}(\lambda)$, as well as the bulk monodromy matrix $T(\lambda)$, act on the tensor product $\mathcal{H}_{0}\otimes\mathcal{H}$, where $\mathcal{H}_{0}\simeq\mathbb{C}^{2}$ is the so-called auxiliary space, whereas $\mathcal{H}=\otimes_{n=1}^{N}\mathcal{H}_{n}$, with $\mathcal{H}_{n}\simeq\mathbb{C}^{2}$, is the quantum space of the model. In other words, $\mathcal{U}_{-}(\lambda)$ and $T(\lambda)$ are to be understood as $\times 2$ 2 matrices (as written in the auxiliary space) with operators entries in $\mathcal{H}$, and the products in (2.14), (2.15) and (2.16) are to be understood as matrix multiplication in $\mathcal{H}_{0}$. In particular, we shall denote
@@ 3.1 Transfer matrix properties
[…]
@@
$\mathcal{T}(\xi_{0}^{(0)})=\tau_{0}\,\mathbb{I}$ $\text{with }\,\tau_{0}=(-1)^{N}\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,{\det}_{q}T(0),$ (3.7)
$\mathcal{T}(\xi_{-1}^{(0)})=\tau_{-1}\,\mathbb{I}$ $\text{with }\,\tau_{-1}=c^{z}_{-}\,c^{z}_{+}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,{\det}_{q}T\Big{(}\frac{\pi}{2}\Big{)},$ (3.8)
$\mathcal{T}(\xi_{-2}^{(0)})=\tau_{-2}\,\mathbb{I}$ $\text{with }\,\tau_{-2}=-e^{i\big{(}2\sum_{n=1}^{N}\xi_{n}-(N+3)\eta-\frac{3\pi\omega}{2}\big{)}}\,c^{x}_{-}\,c_{+}^{x}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,{\det}_{q}T\Big{(}\!-\!\frac{\pi\omega}{2}\Big{)},$ (3.9)
$\mathcal{T}(\xi_{-3}^{(0)})=\tau_{-3}\,\mathbb{I}$ $\text{with }\,\tau_{-3}=e^{i\big{(}2\sum_{n=1}^{N}\xi_{n}-(N+3)\eta-\frac{3\pi\omega}{2}\big{)}}\,c^{y}_{-}\,c_{+}^{y}\,\frac{\operatorname{\theta}(2\eta)}{\operatorname{\theta}(\eta)}\,{\det}_{q}T\Big{(}\!-\!\frac{\pi+\pi\omega}{2}\Big{)}.$ (3.10)
@@ 3.2 SoV basis
Then, one can show similarly as in [1] that, for almost any choice [^5] In [1], the determinant of the matrix formed by the coordinates of the covectors (3.17) in the natural basis is computed and proven nonzero for some special choice of the inhomogeneities and the coordinates of the convector $\langle\,S\,|$. Being this determinant a nonzero multivariable polynomial in these parameters, it can be zero only on the codimension 1 hypersurface of its zeros. This means that these covectors form a basis out of this zero hypersurface, that is for almost any value of the inhomogeneities and choice of $\langle\,S\,|$. of the co-vector $\langle\,S\,|$ and of the inhomogeneity parameters satisfying (3.15), the states
@@ 3.4 Characterisation in terms of a functional $TQ$-equation
[…]
@@
[…]
@@ 4 Scalar products of separate states: main results
for some $\boldsymbol{\varepsilon},\boldsymbol{\varepsilon^{\prime}}\in\{\pm 1\}^{6}$. Here ${\textsc{a}}_{\boldsymbol{\varepsilon}}$ denotes the solution of (3.29) given by (3.32), whereas ${}_{\boldsymbol{\varepsilon}}\langle\,\mathbf{h}\,|$ and $|\,\mathbf{h}\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}$, $\mathbf{h}\equiv(h_{1},\ldots,h_{N})\in\{0,1\}^{N}$, stand for the states (3.17) and (3.20) with function $\mathbf{A}$ (3.16) given by $\mathbf{A}_{\boldsymbol{\varepsilon}}$ and $\mathbf{A}_{\boldsymbol{\varepsilon^{\prime}}}$ respectively777 More generally, the states ${}_{\boldsymbol{\varepsilon}}\langle\,\mathbf{h}\,|$ and $|\,\mathbf{h}\,\rangle_{\boldsymbol{\varepsilon}}$ in (4.1) and (4.4) may stand for any SoV basis elements diagonalising the transfer matrix $\mathcal{T}(\lambda)$ as in Theorem 3.1, provided that it satisfies (3.21) for a given normalisation constant $\mathcal{N}_{0}$, and that it also satisfies (4.3). For instance, such a basis could also be constructed as in [73] by using the Vertex-IRF transformation.. Note that we have
@@ 4.1 The original SoV determinant representation
${}_{\boldsymbol{\varepsilon}}\langle\,P\,|\,Q\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}={}_{\boldsymbol{\varepsilon}}\langle\,Q\,|\,P\,\rangle_{\boldsymbol{\varepsilon^{\prime}}}=\prod_{n=1}^{N}P(\xi_{n}^{(0)})\,Q(\xi_{n}^{(0)})\!$
×$\hskip 56.9055pt\times\sum_{\mathbf{h}\in\{0,1\}^{N}}\prod_{n=1}^{N}\left[-\frac{{\textsc{a}}_{\boldsymbol{\varepsilon}}(\frac{\eta}{2}+\xi_{n})}{{\textsc{a}}_{\boldsymbol{\varepsilon^{\prime}}}(\frac{\eta}{2}-\xi_{n})}\frac{P(\xi_{n}^{(1)})\,Q(\xi_{n}^{(1)})}{P(\xi_{n}^{(0)})\,Q(\xi_{n}^{(0)})}\right]^{h_{n}}\frac{V(\xi_{1}^{(1-h_{1})},\ldots,\xi_{N}^{(1-h_{N})})}{V(\xi_{1},\ldots,\xi_{N})}.$ (4.5)
@@
[…]
@@ 4.2 Slavnov’s type determinant representation
×$\quad\times\bigg{[}-\delta_{j,k}\,\frac{\partial}{\partial\mu}\left(\log\frac{\mathbf{A}_{\boldsymbol{\varepsilon}}(\mu)\,Q(\mu-\eta)}{\mathbf{A}_{\boldsymbol{\varepsilon}}(-\mu)\,Q(\mu+\eta)}\right)_{\mu=q_{i}}+K(q_{i}-q_{j})-K(q_{i}+q_{j})\bigg{]},$ (4.17)
@@ 5 Scalar products of separate states: details of the computations
[…]
@@
[…]
@@
[…]
@@
[…]
@@
To conclude this section, let us briefly comment about the case in which 2 of the 6 boundary parameters are opposite from each others, say for instance $a_{5}=-a_{6}$. We have already mentioned that, in this special case, it was possible to apply the strategy of [4, 5] to compute the scalar products, see the footnote 10. The choice $a_{5}=-a_{6}$ also induces some simplifications in the new approach presented in this section: under this hypothesis, we can indeed simply choose $\gamma=a_{5}$ in (5.12)-(5.1), so that the functions (5.1) no longer have a pole in $\gamma$. This means that the terms of the form $\mathsf{X}_{\gamma}(u)\,\mathsf{M}_{\{a\}}(\gamma,v)$ in the expressions (5.1)-(5.1) vanish. Therefore, in particular, the matrix $\mathcal{G}^{(2,2)}$ (5.1) is simply a diagonal matrix, instead of being the sum of a diagonal matrix and a rank one matrix, so that the explicit expression of its determinant – and hence of the constant $\mathsf{c}_{Q}$ (5.34) – is much simpler. The general formula issued from (5.26), in which we do not suppose $q_{1},\ldots,q_{M}$ to be a solution of the Bethe equations, would also be much less cumbersome in that case. However, this does not induce any particular simplification in our main result (4.14), once we impose the Bethe equations for $q_{1},\ldots,q_{M}$ and normalise adequately the scalar product.
@@
[…]
@@
The last column in the determinant of the numerator of (5.92) being identically 0, we obtain the result.
@@ 6 Conclusion
The new scalar product formulas that we have obtained here constitute the first step towards the computation of the correlation functions. Building on this result, we have implemented the generalisation of the results of [90, 91, 92] to the XYZ open spin chain in [106].
Appendix A The trigonometric limit
We consider here the trigonometric limit $\Im\omega\to+\infty$, in which, if $\lambda$ remains finite,
$\lim_{\Im\omega\rightarrow+\infty}\frac{e^{-i\pi\omega/4}}{2}\,\theta_{1}(\lambda|\omega)=\sin\lambda,\qquad\lim_{\Im\omega\rightarrow+\infty}\frac{e^{-i\pi\omega/4}}{2}\,\theta_{2}(\lambda|\omega)=\cos\lambda,$ (A.1)
$\lim_{\Im\omega\rightarrow+\infty}\theta_{3}(\lambda|\omega)=\lim_{\Im\omega\rightarrow+\infty}\theta_{4}(\lambda|\omega)=1.$ (A.2)
Let us define new boundary parameters $\varphi_{\pm},\psi_{\pm}$ and $\tau_{\pm}$ by
$\alpha_{1}^{\pm}=-i\varphi_{\mp},\qquad\alpha_{2}^{\pm}=i\psi_{\mp}-\epsilon\frac{\pi}{2},\qquad\alpha_{3}^{\pm}=i\tau_{\mp}+\epsilon\frac{\pi}{2}+\frac{\pi\omega}{2},$ (A.3)
for some arbitrary sign $\epsilon$, and we suppose that these new boundary parameters $\varphi_{\pm},\psi_{\pm},\tau_{\pm}$ remain finite in the limit $\Im\omega\to+\infty$. Then, the boundary parameters $c^{x,y,z}_{\pm}$ (2.9) behave in this limit in terms of (A.3) as
$\lim_{\Im\omega\rightarrow+\infty}2e^{i\pi\omega/4}c^{x}_{\mp}=\frac{-i\cosh(\tau^{\pm})}{\sinh(\varphi^{\pm})\,\cosh(\psi^{\pm})},$ (A.4)
$\lim_{\Im\omega\rightarrow+\infty}2e^{i\pi\omega/4}c^{y}_{\mp}=\frac{-i\sinh(\tau^{\pm})}{\sinh(\varphi^{\pm})\,\cosh(\psi^{\pm})},$ (A.5)
$\lim_{\Im\omega\rightarrow+\infty}c^{z}_{\mp}=-i\frac{\cosh(\varphi^{\pm})\,\sinh(\psi^{\pm})}{\sinh(\varphi^{\pm})\,\cosh(\psi^{\pm})}.$ (A.6)
Hence, in the limit $\Im\omega\to+\infty$, with $\eta,\varphi_{\pm},\psi_{\pm},\tau_{\pm}$ remaining finite, the Hamiltonian (2.27) degenerates into the Hamiltonian of the open XXZ spin chain:
$H_{\text{XXZ}}=\sum_{n=1}^{N-1}\Big{[}\sigma_{n}^{x}\sigma_{n+1}^{x}+\sigma_{n}^{y}\sigma_{n+1}^{y}+\Delta\,\sigma_{n}^{z}\sigma_{n+1}^{z}\Big{]}+\sum_{a\in\{x,y,z\}}\Big{[}\tilde{h}_{-}^{a}\,\sigma_{1}^{a}+\tilde{h}_{+}^{a}\sigma_{N}^{a}\Big{]},$ (A.7)
in which
$\Delta=\cos\eta=\cosh\tilde{\eta},\qquad\text{with}\quad\eta=i\tilde{\eta},$ (A.8)
$\tilde{h}_{\pm}^{x}=\sinh\tilde{\eta}\,\frac{\cosh\tau_{\pm}}{\sinh\varphi_{\pm}\,\cosh\psi_{\pm}},$ (A.9)
$\tilde{h}_{\pm}^{y}=i\sinh\tilde{\eta}\,\frac{\sinh\tau_{\pm}}{\sinh\varphi_{\pm}\,\cosh\psi_{\pm}},$ (A.10)
$\tilde{h}_{\pm}^{z}=\sinh\tilde{\eta}\,\coth\varphi_{\pm}\,\tanh\psi_{\pm}.$ (A.11)
Moreover, still in this limit, if in addition $\lambda=iu$ remains finite, the R-matrix (2.1) degenerates, up to a global normalisation, into the 6-vertex R-matrix,
$R_{\text{XXZ}}(u)=\begin{pmatrix}\sinh(u+\tilde{\eta})&0&0&0\\ 0&\sinh u&\sinh\tilde{\eta}&0\\ 0&\sinh\tilde{\eta}&\sinh u&0\\ 0&0&0&\sinh(u+\tilde{\eta})\end{pmatrix},$ (A.12)
and the boundary K-matrices (2.13) become
$K_{\pm}(\lambda\mp\eta/2)$ $\underset{\Im\omega\to+\infty}{\longrightarrow}K_{\text{XXZ}}(-i\lambda;\varphi_{\mp},\psi_{\mp},\tau_{\mp}),$ (A.13)
in which
$K_{\text{XXZ}}(u;\varphi,\psi,\tau)$ $=\cosh u\,\mathbb{I}+\frac{\cosh\varphi\,\sinh\psi}{\sinh\varphi\,\cosh\psi}\,\sinh u\,\sigma^{z}+\frac{\sinh(2u)}{2\sinh\varphi\,\cosh\psi}\,\left[\cosh\tau\,\sigma^{x}+i\sinh\tau\,\sigma^{y}\right]$
$=\begin{pmatrix}\cosh u+\sinh u\frac{\cosh\varphi\sinh\psi}{\sinh\varphi\cosh\psi}&e^{\tau}\frac{\sinh 2u}{2\sinh\varphi\cosh\psi}\\ e^{-\tau}\frac{\sinh 2u}{2\sinh\varphi\cosh\psi}&\cosh u-\sinh u\frac{\cosh\varphi\sinh\psi}{\sinh\varphi\cosh\psi}\end{pmatrix}$ (A.14)
is the K-matrix of the open XXZ spin chain. Finally, the coefficient (3.32) becomes
${\textsc{a}}_{\boldsymbol{\varepsilon}}(\lambda)$ $\underset{\Im\omega\to+\infty}{\longrightarrow}e^{-(\epsilon_{3}^{+}+\epsilon_{3}^{-})(u-\frac{\tilde{\eta}}{2})}\prod_{\sigma=\pm}\frac{\sinh(u-\frac{\tilde{\eta}}{2}+\epsilon_{1}^{\sigma}\varphi_{\sigma})}{\sinh(\epsilon_{1}^{\sigma}\varphi_{\sigma})}\frac{\cosh(u-\frac{\tilde{\eta}}{2}-\epsilon_{2}^{\sigma}\psi_{\sigma})}{\cosh(-\epsilon_{2}^{\sigma}\psi_{\sigma})}.$ (A.15)
For the particular choice $\epsilon_{3}^{+}=-\epsilon_{3}^{-}$, $\epsilon_{1}^{\pm}=\epsilon_{\varphi^{\mp}}$, $\epsilon_{2}^{\pm}=\pm\epsilon_{\psi^{\mp}}$, we recover the convention of our previous XXZ papers [5, 92] [^12] This corresponds more precisely to the convention of [92]. In [5], the notation $\alpha_{\pm},\beta_{\pm}$ was used instead of $\varphi_{\mp},\psi_{\mp}$ respectively.: the constraint (3.35) becomes
$\epsilon_{3}^{+}(\tau_{+}-\tau_{-})+\sum_{\sigma=\pm}\left[\epsilon_{\varphi_{\sigma}}\varphi_{\sigma}+\sigma\epsilon_{\psi_{\sigma}}(\psi_{\sigma}+i\epsilon\frac{\pi}{2})\right]+(N-2M-1)\tilde{\eta}=0,$ (A.16)
with $\epsilon_{\varphi_{+}}\epsilon_{\varphi_{-}}\epsilon_{\psi_{+}}\epsilon_{\psi_{-}}=1$; under this constraint, we recover the XXZ analog of Proposition 3.2 for trigonometric polynomials $Q$ of the form
$Q(u)=\prod_{n=1}^{M}\sinh(u-\tilde{q}_{n})\sinh(u+\tilde{q}_{n}),\qquad\mathbf{\tilde{q}}=(\tilde{q}_{1},\ldots,\tilde{q}_{M})\in\mathbb{C}^{M}.$ (A.17)
It is now easy to see that Theorem 4.1 reduces to Theorem 5.2 of [5] in the trigonometric limit [^13] Up to the normalisation that we have not computed here, since, as already mentioned, it has a complicated expression and is not relevant for the computation of correlation functions..